The sum of the ages of a boy and his mother is 48. If the mother is three more than twice the
boy’s age, find the difference in the ages of the boy and his mother.

Answer by aliokur(1)

B=Ages of boy
M=Ages of mother
B+M=48 => M=48-B and B=48-M
2B+3=M
2B+3=M => 2B+3=48-B => 3B=51 => B=17
M=48-B=48-17=31
M-B=31-17=14